OpenStudy (anonymous):

Philippa has a gateleg table in her kitchen. The leg swings from one corner of the table to the middle of the far side of the table top, as shown in the sketch below.

OpenStudy (anonymous):



OpenStudy (anonymous):

How long is the gateleg of the table?

OpenStudy (anonymous):

@hoblos sorry i need your help again. Pythagorean Theorem again

OpenStudy (hoblos):

 can you apply the formula here ?

OpenStudy (anonymous):

would it be \[g ^{2}=45^{2}+56^{2}?\]

OpenStudy (hoblos):

it's (56/2)^2

OpenStudy (anonymous):

\[(56\div2)^{2}???\]

OpenStudy (hoblos):

\[g^{2} = 45^{2} + (\frac{ 56 }{ 2 })^{2}\]

OpenStudy (hoblos):

yes

OpenStudy (anonymous):

\[56\div2=28^{2}=28\times28=784?\]

OpenStudy (anonymous):

\[g ^{2}=45^{2}+784?\]

OpenStudy (anonymous):

\[g ^{2}=2025+784?\]

OpenStudy (anonymous):

\[g ^{2}=\sqrt{2809}\]

OpenStudy (anonymous):

\[g=53\]

OpenStudy (hoblos):

sorry i couldn't reply yesterday, you got the right answer though good work :D